OpenStudy (anonymous):

A mother wants to give her two children curently aged 8 and 5 money for their birthdays. She wants to give them each R5000 at 13, R7000 at 16, and R10 000 at 18. The bank is offering her interst of 11,2% p.a compounded monthly. What must she invest now so that the children can get teir future birthday presents.

OpenStudy (callisto):

Does she invest every year, or just a single year?

OpenStudy (anonymous):

She invests just once

OpenStudy (callisto):

Use the formula: FV = PV (1+i%)^n for the child aged 8 now let x1 be the amount she needs to invest to get 5000 5000 = PV [1+(11.2%/12)]^ [(13-8)x12] x1 = 5000 / [1+(11.2%/12)]^ [(13-8)x12] let x2 be the amount she needs to invest to get 7000 7000 = PV [1+(11.2%/12)]^ [(16-8)x12] x2 = 7000 / [1+(11.2%/12)]^ [(16-8)x12] let x3 be the amount she needs to invest to get 10000 10000 = PV [1+(11.2%/12)]^ [(18-8)x12] x3 = 10000/ [1+(11.2%/12)]^ [(18-8)x12] Sum of money she needs to invest for her child aged 8 = x1 + x2+x3 For the second child, just do the same Total money she needs to invest = amount she invest for first child + amount she invest for second child Can you get the answer?

OpenStudy (anonymous):

Tanx Callisto

OpenStudy (callisto):

welcome :)